OpenStudy (anonymous):

Your friend hands you a graph of the engine power statistics of a race car. He says "I know this is a translated graph of either x2 or x3, but I can't figure out which one." Using complete sentences, explain to your friend how to tell if it is an even or odd function, and how to determine the amount the graph has been translated. ill give out medals and i just need someone to see if my respond i s correct

OpenStudy (anonymous):

@PinkSapphire

OpenStudy (anonymous):

For x^2, both ends would be either going towards the negative or the positive , which indicates that this function is even. For x^3, one end would be going towards negative, and the other end would be going towards the positive, which indicates that this function is odd. once the basic information has be informed, now the only thing to do is compare the graph with the plain of x^2 and x^3 and apply translation rules to determine the translation

OpenStudy (anonymous):

thats my respond

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@TammisaurusRex

OpenStudy (pinksapphire):

I think that's correct.

OpenStudy (anonymous):

okay im concerned about a couple of more i did can you correct it pleasee

OpenStudy (pinksapphire):

Sure.

OpenStudy (anonymous):

You are having a meeting with the CEO of a technology company. You have interpreted the number of laptops produced versus profit as the function P(x) = x4 -3x3 -8x2 + 12x + 16. Describe to the CEO what the graph looks like. Use complete sentences and focus on the end behaviors of the graph and where the company will break even (where P(x) = 0).

OpenStudy (anonymous):

my response:

OpenStudy (anonymous):

For this presentation you can basically look where the highest point is in the function which is x^4, the rest of the equations are unnecessary to be there, its almost like a distraction, which is known as the end power. Function: P(x) = x^4 -3x^3 -8x^2 + 12x + 16 To predict the end behavior of a polynomial it has to be positive or negative, check if the function is a odd degree or a even degree function and also see if the leading coefficient is positive or negative. since the function is mainly positive the end behavior would be: f(x)(+infinite)-->(-xinfinite), f(x)(+infinite)-->(+infinite) basically the graph wouldn't be v-shaped parabola. it would mostly be wave like lines with the endpoints facing up.

OpenStudy (pinksapphire):

It's a wave of lines.

OpenStudy (anonymous):

i was describing what the parabola would look like

OpenStudy (pinksapphire):

Yes you are right.

OpenStudy (anonymous):

okay a couple more.

OpenStudy (pinksapphire):

Okay...

OpenStudy (anonymous):

Two students in your class, Hunter and Maggie, are disputing a function. Hunter says that for the function, between x = -2 and x = 2, the average rate of change is 0. Maggie says that for the function, between x = -2 and x = 2, the graph goes up through a turning point, and then back down. Explain how Hunter and Maggie can both be correct, using complete sentences.

OpenStudy (anonymous):

Maggie is correct because she said " the graph goes up through the turning point, and then back down". explanation: when both both ends in the same direction, it goes up and down. therefore it would be an even function and is x= -2, x= 2. Hunter is incorrect. explanation: the average rate consists of dividing delta y and delta x. which would look like this: (y1-y2)/(x1-x2) as you can see there's not enough of information on his part, and even if he used another kind of function to solve this problem the answer wouldn't be 0 it would be 1 or -1

OpenStudy (anonymous):

@PinkSapphire

OpenStudy (pinksapphire):

I think you're right on this one but I'm not 100% sure...

OpenStudy (anonymous):

okay thanks now only 2 more 1 of the questions i guessed on.

OpenStudy (pinksapphire):

Okay.

OpenStudy (anonymous):

this is the 1 i guessed on

OpenStudy (anonymous):

The change in water level of a lake is modeled by a polynomial function, W(x). Describe how to find the x-intercepts of W(x) and how to construct a rough graph of W(x) so that the Parks Department can predict when there will be no change in the water level. You may create a sample polynomial to be used in your explanations.

OpenStudy (anonymous):

i made this up: f(x)= f(x)= x^0+2x^2-x-2 the possible zeros are going to be +-1, +-2 set it to equal zero: 0=(1)^3+2(1)^1-(1) --> 1+2-1-2=0 so basically 1 works for this problem. 0= x^0+2x^2-x-2 0=(-1)^2+2(-1)^2-1(-1)-2 --> -1+2+1-2=0 so basically -1 works for this problem. these are the x-intercepts 1,-1&-2,or(1,0)(-1,0)

OpenStudy (anonymous):

@PinkSapphire

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (pinksapphire):

I don't know about this one...

OpenStudy (anonymous):

me either okay the last 1

OpenStudy (anonymous):

In your lab, a substance's temperature has been observed to follow the function T(x) = (x + 5)3 + 7. The turning point of the graph is where the substance changes from a liquid to a solid. Explain to your fellow scientists how to find the turning point of this function, using complete sentences.

OpenStudy (anonymous):

first off we have to take the derivative of t(x) setting it to 0, and solving for 0. when we have solved for the equation we would get -5 so it would be x= -5. since we have a double solution at x= -5 there would be an inflection point. basically what that means is at the turning point of the graph is where the changes from liquid to solid.

OpenStudy (anonymous):

@PinkSapphire

OpenStudy (pinksapphire):

How did you get -5?

OpenStudy (anonymous):

when i five into x i ended up getting negative 5

OpenStudy (anonymous):

is my explanation wrong

OpenStudy (anonymous):

@PinkSapphire

OpenStudy (pinksapphire):

I don't know...I don't want to give you the wrong answer so can you ask someone else? I'm so sorry :( You should ask @Hero.

OpenStudy (anonymous):

okay thankz :)

OpenStudy (anonymous):

@beccaboo333

OpenStudy (beccaboo333):

I'd help if I knew how to math. I'm sorry math is my least subject.

OpenStudy (anonymous):

okay its fine

OpenStudy (anonymous):

@ParthKohli